OpenStudy (anonymous):

what multiplys to get 18 and adds to get 3?

OpenStudy (saifoo.khan):

18x+y = 3

OpenStudy (saifoo.khan):

o_O

OpenStudy (bahrom7893):

x*y = 18 x+y = 3 => x = 3-y (3-y)*y = 18 3y - y^2 - 18 = 0

OpenStudy (bahrom7893):

just solve the equations i gotta go help someone